OpenStudy (anonymous):

HELP PLEASE Two lines, A and B, are represented by the equations given below: Line A: x + y = 6 Line B: x + y = 4 Which statement is true about the solution to the set of equations? There are infinitely many solutions. There is no solution. It is (6, 4). It is (4, 6).

OpenStudy (anonymous):

Please HELP

Mehek (mehek14):

what did you choose?

OpenStudy (anonymous):

D

Mehek (mehek14):

that is incorrect

OpenStudy (anonymous):

So which one is it then

OpenStudy (anonymous):

Or how do I find the correct one

Mehek (mehek14):

change x + y = 6 into y = -x + 6 and x + y = 4 into y = -x + 4 they are almost the same except the y-intercepts are different that means they are parallel lines so there is no solution since parallel lines never intersect

OpenStudy (anonymous):

oh

Mehek (mehek14):

so B

Mehek (mehek14):

http://prntscr.com/7sreg6